Andy's average driving speed for a 4-hour trip was 45 mph. During that first 3 hours he drove 40 mph. What was his average speed for the last hour of his trip?
Answer by Tieria.Erde(4)

Calculate the total distance the four hour trip was. 1) [4*45=180]
Then calculate the first 3 hours. 2) [3*40=120]
Subtract the difference [180-120=60]
The answer is 60mph